OpenStudy (anonymous):

Find the value of x and y in the parallelogram below one moment im drawing it o.o/

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

in any parallelogram, the opposite sides are congruent or equal in length

jimthompson5910 (jim_thompson5910):

so AB = CD 3x - 6 = 2x+8 solve for x

jimthompson5910 (jim_thompson5910):

also, in any parallelogram, the adjacent angles are supplementary so they add to 180 degrees (Angle A) + (angle D) = 180 (64) + (4y - 2) = 180 solve for y

OpenStudy (anonymous):

does x=14? im having trouble with solving for y i think i do: 64+4y-2=180 64-2+4y=180 62+4y=180 -62 -62 ------------------- 4y=118 ---------- 4 y=29.5?

jimthompson5910 (jim_thompson5910):

both x and y are correct, nice work

jimthompson5910 (jim_thompson5910):

optionally you can have y as a fraction and have y = 59/2, but y = 29.5 works too

OpenStudy (anonymous):

thank you so much!:D

jimthompson5910 (jim_thompson5910):

you're welcome